OpenStudy (anonymous):

What are the solutions to the following system of equations? y = x2 + 12x + 30 8x − y = 10 (−4, −2) and (2, 5) (−2, −4) and (2, 5) (−2, −4) and (5, 2) No Real Solutions

OpenStudy (anonymous):

@ivettef365 please helppp

OpenStudy (anonymous):

hope this helps :)) http://openstudy.com/study#/updates/5217503be4b00ebc7dc0e8bf

OpenStudy (anonymous):

maybe this one will be better :) http://openstudy.com/study#/updates/52174c66e4b00ebc7dc0e810

OpenStudy (ybarrap):

$$ \tt \bf\\ \Large y = x^2 + 12x + 30\\ \Large 8x - y = 10 \implies y=8x-10\\ \Large \text{Set equal:}\\ \Large x^2+12x+30=8x-10\\ \Large x^2+4x+40=0\\ \Large (x+2)^2+40-4=0\\ \Large (x+2)^2=-36\\ \Large x+2=\pm i\sqrt{36}\\ \Large x=-2\pm i\sqrt{36}\\ \Large \text{roots are complex} $$